# ESC: Efficient Speech Coding with Cross-Scale Residual Vector Quantized Transformers

### v2

@@ Abstract
Neural speech codecs aim to compress input signals into minimal bits while maintaining content quality in a low-latency manner. However, existing codecs often trade computational complexity for reconstruction performance. These codecs primarily use convolutional blocks for feature transformation layers, which are not inherently suited for capturing the local redundancies in speech signals. To compensate, they require either adversarial discriminators or a large number of model parameters to enhance audio quality. In response to these challenges, we introduce the Efficient Speech Codec (ESC), a lightweight, parameter-efficient speech codec based on a cross-scale residual vector quantization scheme and transformers. Our model employs mirrored hierarchical window transformer blocks and performs step-wise decoding from coarse-to-fine feature representations. To enhance bitrate efficiency, we propose a novel combination of vector quantization techniques along with a pre-training paradigm. Extensive experiments demonstrate that ESC can achieve high-fidelity speech reconstruction with significantly lower complexity, making it a promising alternative to existing convolutional audio codecs [^1] Our code and pretrained models are available here..
@@ 1 Introduction
Recent advancements in deep learning have demonstrated the superiority of neural speech codecs over traditional ones, which rely on complex expert design and psycho-acoustic knowledge Valin et al. (2012); Dietz et al. (2015). Early efforts integrating deep generative models such as WaveNet Oord et al. (2016) and SampleRNN Mehri et al. (2017) into audio codecs have yielded promising results. These models serve as powerful decoders, synthesizing high-quality speech audio based on the intermediate representations generated by traditional codecs Kleijn et al. (2018); Klejsa et al. (2019). However, their auto-regressive decoding nature often introduces significant inference latency, limiting their practical application.
@@
Alternatively, some end-to-end neural audio codecs build upon the vector quantization (VQ) network first introduced by Van Den Oord et al. (2017). VQ networks, which consist of trainable codebooks, can discretize continuous input vectors by assigning them to the nearest neighboring code-vectors, making them suitable for both generation and compression tasks. Based on this framework, existing VQ codecs Zeghidour et al. (2021); Défossez et al. (2023); Kumar et al. (2023) typically follow a three-stage architecture with a convolutional encoder and decoder interleaved with a residual vector quantization (RVQ) module Vasuki and Vanathi (2006) at the bottleneck. The encoder and decoder perform downsampling and upsampling of audio signals in the time-domain to produce hierarchical representations. RVQ is an enhanced quantization scheme where a series of VQs recursively quantize the residuals from previous stages. Additionally, these codecs utilize generative adversarial networks (GANs) to remove artifacts and produce high-fidelity audio reconstructions. Significant effort has been dedicated to developing effective audio discriminators that leverage multi-scale features of waveforms and spectrograms to improve the generated audio during training Esser et al. (2021); Kumar et al. (2019); Kong et al. (2020). Currently, VQ-based codecs have demonstrated remarkable capabilities in audio reconstruction even at ultra-low bitrates.
@@
Despite these advantages, we find that convolutional VQ-based codecs heavily rely on powerful GAN discriminators to produce high-quality audio, introducing additional optimization challenges due to adversarial training. Moreover, convolutional codecs tend to face computational limitations, as they require a large number of parameters to achieve a good trade-off between compression rate and reconstruction quality. To address these issues, our work aims to develop a more parameter-efficient speech codec by implementing the following architectural improvements: 1) Replacing convolutional layers with efficient Swin-Transformer Blocks (STBs) Liu et al. (2021), which have greater capacity for modeling acoustic features; 2) Utilizing the cross-scale residual vector quantization (CS-RVQ) scheme Jiang et al. (2022a) instead of RVQ, extending quantization from a fixed level to multiple levels.
@@
In addition, training VQ-based codecs encounters a significant obstacle: a fraction of the vector quantizer’s codebook is often underutilized for representing input vectors, leading to a phenomenon known as codebook collapse. This issue is frequently observed in image generation tasks Takida et al. (2022); Zhang et al. (2023); Huh et al. (2023). To address this in the context of speech compression, we propose combining product vector quantization (PVQ) Baevski et al. (2019), latent dimension reduction Yu et al. (2022), and Euclidean normalization Łańcucki et al. (2020) to enhance codebook utilization. Additionally, we introduce a learning paradigm to facilate codebook optimization. It incorporates a pre-training stage during which VQs are freezed and trained afterwards.
In summary, the contributions of our work are as follows:
  • We introduce ESC, a lightweight efficient speech codec that integrates cross-scale residual vector quantization (CS-RVQ) and mirrored hierarchical Swin-Transformer blocks, achieving a superior trade-off between reconstruction performance and codec complexity compared to current state-of-the-art models.
@@
  • We propose a novel combination of vector quantization techniques along with a pre-training paradigm that includes a warm start for the encoder and decoder, aimed at mitigating codebook collapse and enhancing bitrate efficiency.
@@
  • Our extensive experiments demonstrate that transformers and CS-RVQ, the core components of ESC, are more efficient backbones for speech compression than convolutions and RVQ.
@@ 2 Related Work
2.1 VQ-based Audio Codecs
Notable examples in recent VQ-based audio codecs include SoundStream Zeghidour et al. (2021), EnCodec Défossez et al. (2023), and Descript’s Audio Codec (DAC) Kumar et al. (2023). SoundStream is recognized as the first universal neural codec designed for various audio types. EnCodec enhances compression rates by integrating a lightweight transformer language model at the discrete latent space and explores a streaming architecture. Following similar methodologies, Kumar et al. (2023) identify the side effects of quantization dropout, a technique for bitrate scalability, and demonstrate the superiority of a periodic inductive bias function over other common activation functions in audio signal modeling gil Lee et al. (2023); Ziyin et al. (2020). These codecs process audio signals directly in the time domain and are thus categorized as time-domain codecs. Meanwhile, frequency-domain codecs, such as TFNet Jiang et al. (2022b) and our proposed ESC, leverage the invertible short-time Fourier transform (STFT) to convert time-domain audio into spectrograms. The encoder and decoder thus process these two-dimensional spectral features directly. Among recent neural codecs, DAC achieves state-of-the-art compression ratios and audio reconstruction quality, although its computation bottleneck is ignored.
@@ 2.2 Swin Transformers
Vision Transformers (ViTs) Dosovitskiy et al. (2020) have outperformed convolutional neural networks (CNNs) in various image processing tasks due to their superior ability to capture complex patterns. The Swin Transformer Liu et al. (2021), a notable variant, enhances this capability through a hierarchical approach that leverages window attention mechanisms. It can scale efficiently to high-resolution signals while maintaining computational efficiency. In image compression, Swin Transformers have demonstrated exceptional performance. Studies by Zhu et al. (2021) and Zou et al. (2022) have shown that they outperform CNNs in modeling spatial hierarchies and long-range dependencies. The attention mechanisms enable the accurate preservation of essential details and textures, even at lower bitrates. These capabilities highlight the potential of transformers for applications beyond image compression, suggesting their suitability for the field of speech compression as well.
@@ 2.3 Vector Quantization
Proposed in Vector Quantized Variational Autoencoders (VQ-VAEs) (Van Den Oord et al., 2017), vector quantization (VQ) serves as a trainable layer that deterministically quantizes the encoded latent variables to their nearest neighbors in an embedding codebook. A vector quantization layer $Q(\cdot\ ;\mathcal{C})$ is parameterized by a codebook collection of continuous vectors $\mathcal{C}={\bm{c}_{1},...,\bm{c}_{K}}$. Each vector is referred to as a codeword, and its index is known as a code. The layer quantizes a vector $\bm{z}_{e}\in\mathbb{R}^{d}$ to be $\bm{z}_{q}\in\mathbb{R}^{d}$ by selecting the Euclidean nearest codeword $\bm{c}_{k}$ from the codebook $\mathcal{C}$, i.e.,
@@
For notation convenience, we denote the vector quantization function as $\bm{z}_{e}\xrightarrow[]{Q(\cdot;\mathcal{C})}(\bm{z}_{q},\tilde{z}_{q})$, where $\tilde{z}_{q}:=k$ is the corresponding discrete code for the nearest codeword. During compression, the encoding process outputs the discrete index $\tilde{z}_{q}$, which is stored with a $\log_{2}K$ bit budget. The decoding process starts by retrieving the continuous vector $\bm{z}_{q}$ from the codebook using the encoded index $\tilde{z}_{q}$. The VQ function $Q(\cdot;\mathcal{C})$ is non-differentiable due to the $\operatorname*{arg\,min}$ operator. Common strategies use a straight-through estimator (STE) Bengio et al. (2013) to bypass this in back-propagation. In other words, the gradient component $\frac{\partial\bm{z}_{q}}{\partial\bm{z}e}$ is eliminated, and as compensation, a codebook loss and a commitment loss are proposed to pull together the codewords and the encoded latents:
@@
Here $\text{sg}(\cdot)$ denotes the stop gradient operator. The first term updates the codebook with an $l_{2}$ error, moving the codewords towards the encoder outputs. The second term ensures $\bm{z}_{e}$ can commit to the embedding without growing arbitrarily. The scalar $\beta$ is used to balance the importance of updating the codebook and encoder.
@@ 2.4 Codebook Collapse
Straight-through estimators (STEs) often lead to significant issues, most notably the issue of codebook collapse Vuong et al. (2023). In a recent study, Huh et al. (2023) offer a straightforward explanation, attributing it to an internal codebook covariate shift during the training phase. Frequent adjustments in encoder representations cause misalignment with VQ codebooks, resulting in only a subset of codewords being updated. Consequently, VQ layers are prone to divergence, often ending up with a significant number of dead vectors. Various strategies have been proposed in generative modeling tasks to address this problem, including the use of stochastic VQs Takida et al. (2022); Zhang et al. (2023), self-annealing soft-to-hard quantization approaches Agustsson et al. (2017), re-initializing codewords with K-means centroids every few iterations Łańcucki et al. (2020); Dhariwal et al. (2020), and avoiding the issues of STEs with finite scalar quantizers Mentzer et al. (2024). In audio compression, Kumar et al. (2023) attempt to mitigate codebook collapse by reducing the dimensions of codewords Yu et al. (2022) and normalizing them within a Euclidean ball Łańcucki et al. (2020).
Figure 1: The framework of ESC: Input speech is transformed to a complex STFT $\mathcal{X}$ and linearly embedded into patches. Encoder STBs iteratively halve the frequency resolution and produce hierarchical feature representations. Mirrored decoder STBs iteratively recover the frequency resolution by leveraging coarse-to-fine encoded features. The intermediate residuals between encoder and decoder hidden states are progressively quantized to refine decoding. The entire network is solely composed of efficient transformer blocks. The figure displays a scenario when the deepest $3$ of $n+1$ total bitstreams (solid lines) are transmitted, with others left inactive.
@@ 3.1 Overall Architecture
As illustrated in Figure 1, our codec operates on the complex spectrum $\mathcal{X}\in\mathbb{R}^{2\times F\times T}$ derived from the Short-Time Fourier Transform (STFT) of a speech signal. Here, the real and imaginary components of $\mathcal{X}$ are treated as separate channels. The model consists of a series of mirrored transformer encoder and decoder layers, each performing downsampling or upsampling to generate coarse and fine representations, as detailed in Section 3.3. Starting from the quantized latents at the bottleneck VQ, the decoder progressively reconstructs the original spectrum by leveraging multi-level residuals between the intermediate features of the encoder and decoder. This cross-scale decoding mechanism is explained in further detail in Section 3.4. Finally, the reconstructed spectrum $\hat{\mathcal{X}}$ is transformed back into the time domain through the inverse STFT.
@@ 3.2 Notations
We first define some architectural notations for clarity. We represent the encoder and decoder functions $F_{\phi}(\cdot)$ and $G_{\psi}(\cdot)$, each as a composition of individual layer functions $f_{\phi_{1}},\ldots,f_{\phi_{n}}$ and $g_{\psi_{1}},\ldots,g_{\psi_{n}}$, respectively. We use $\mathcal{Z}$ to represent spectrum feature and $\bm{z}$ to denote a feature frame. Specifically, we denote $\mathcal{Z}_{e_{i}}$ as the hidden feature after $i$ encoder layers, and $\mathcal{Z}_{q_{i}}$ as the corresponding $i$-th decoder feature. Formally, this is expressed as:
@@
where $\mathcal{Z}_{e_{0}}$ is the original input feature and $\mathcal{Z}_{q_{0}}$ is the latent representation at the bottleneck.
@@ 3.3 Transformer Encoder and Decoder
Patchify. The encoder begins with a linear patchify module, where the complex spectrum $\mathcal{X}$ is divided into small patches and linearly projected:
@@
Here, the patch size across the frequency and temporal dimensions is $(\frac{F}{H_{0}},\frac{T}{W_{0}})$. This step reduces the input resolution to alleviate the computational burden on transformers. At the end of the decoder, a symmetric de-patchify module reshapes the decoded patch embedding $\mathcal{Z}_{q_{n}}$ and linearly projects it back to produce a recovered spectrum $\hat{\mathcal{X}}$.
@@
Downsampling and upsampling. Our codec maintains temporal resolution while scaling frequency resolution to equalize bitrates across different bitstreams. To achieve this, we modify the original patch merging/splitting modules with a single-dimensional pixel unshuffle/shuffle module Shi et al. (2016) along the frequency dimension. During encoder downsampling, an intermediate encoder spectrum feature $\mathcal{Z}_{e_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is first reshaped and then projected as follows:
@@
where $v$ is the scaling factor. The upsampling process mirrors this operation in reverse. An intermediate decoder feature $\mathcal{Z}_{q_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is up-projected and then reshaped, resulting in an up-scaled frequency resolution:
@@
Thus, the transformer encoder and decoder layers are mirrored, creating symmetric and hierarchical representations of the input audio spectrum. With these backbones, ESC is a fully transformer-based codec without any convolutional modules.

### v3

@@ Abstract
Neural speech codecs aim to compress input signals into minimal bits while maintaining content quality in a low-latency manner. However, existing neural codecs often trade model complexity for reconstruction performance. These codecs primarily use convolutional blocks for feature transformation, which are not inherently suited for capturing the local redundancies in speech signals. To compensate, they require either adversarial discriminators or a large number of model parameters to enhance audio quality. In response to these challenges, we introduce the Efficient Speech Codec (ESC) [^1] Code and pretrained models available at https://github.com/yzGuu830/efficient-speech-codec, a lightweight, parameter-efficient speech codec based on a cross-scale residual vector quantization scheme and transformers. Our model employs mirrored hierarchical window transformer blocks and performs step-wise decoding from coarse-to-fine feature representations. To enhance bitrate efficiency, we propose a novel combination of vector quantization techniques along with a pre-training paradigm. Extensive experiments demonstrate that ESC can achieve high-fidelity speech reconstruction with significantly lower model complexity, making it a promising alternative to existing convolutional audio codecs.
@@ 1 Introduction
Recent advancements in deep learning have demonstrated the superiority of neural speech codecs over traditional ones, which rely on complex expert design and psycho-acoustic knowledge Valin et al. (2012); Dietz et al. (2015). Early efforts integrating deep generative models, such as WaveNet Oord et al. (2016) and SampleRNN Mehri et al. (2017), into audio codecs have delivered promising results. These models, acting as powerful decoders, synthesize high-quality speech from intermediate representations produced by traditional codecs Kleijn et al. (2018); Klejsa et al. (2019). However, their auto-regressive nature of the decoding process often introduces significant inference latency, limiting their practical application.
@@
Alternatively, some end-to-end neural audio codecs leverage the vector quantization (VQ) network first introduced by Van Den Oord et al. (2017). VQ networks use a learnable collection of code-vectors, known as a codebook, to quantize continuous vectors by assigning them to the nearest codeword. This discretization positions VQNs well-suited for both generation and compression tasks. Following this approach, existing VQ codecs Zeghidour et al. (2021); Défossez et al. (2023); Kumar et al. (2023) typically employ a three-stage architecture: a convolutional encoder and decoder, and a residual vector quantization (RVQ) module Vasuki and Vanathi (2006) applied in the latent space. The encoder and decoder downsample and upsample audio waveform features, creating hierarchical representations. RVQ further refines vanilla vector quantization by minimizing quantization error through a series of codebooks that recursively quantize the residuals from previous stages. Additionally, these codecs employ adversarial discriminators to remove artifacts and produce high-fidelity audio reconstructions. Substantial effort has been dedicated to designing effective audio discriminators, including an improved feature matching loss Kumar et al. (2019), as well as various multi-resolution waveform and spectrogram discriminators Kong et al. (2020); Zeghidour et al. (2021); Défossez et al. (2023); gil Lee et al. (2023). VQ-based audio codecs have demonstrated remarkable performance in audio reconstruction, even at ultra-low bitrates.
@@
Despite these advantages, we find that convolutional VQ codecs heavily depend on powerful discriminators to produce high-quality audio, posing additional optimization challenges due to adversarial training. Moreover, these codecs tend to confront computational constraints, as they require a large number of parameters to balance high compression rates and reconstruction performance. To address these issues, our work develops a more parameter-efficient speech codec by reducing model complexity and implementing the following architectural improvements: 1) replacing convolutional layers with efficient Swin-Transformer Blocks (STBs) Liu et al. (2021), which can better model acoustic features; 2) utilizing the cross-scale residual vector quantization (CS-RVQ) scheme Jiang et al. (2022a) instead of RVQ, extending quantization from a fixed level to multiple levels.
@@
In addition, training VQ codecs frequently leads to a significant challenge: codebook collapse, where a fraction of the codebook remains underutilized in representing input vectors. This issue is frequently observed when training visual tokenizers for generative vision tasks Takida et al. (2022); Zhang et al. (2023); Huh et al. (2023). To address this problem in speech compression, we propose combining product vector quantization (PVQ) Baevski et al. (2019), code factorization Yu et al. (2022), and Euclidean normalization Łańcucki et al. (2020) to enhance codebook utilization. Furthermore, we introduce a learning paradigm to facilitate optimization, which includes a pre-training stage where the codebooks are deactivated and trained subsequently.
In summary, the key contributions of our work are as follows:
  • We introduce ESC, a fully transformer-based speech codec with cross-scale quantization structures. It achieves a superior tradeoff between compression rate, reconstruction quality, and model complexity, outperforming current state-of-the-art models.
@@
  • We propose a novel combination of vector quantization techniques within the cross-scale residual vector quantization (CS-RVQ) framework, coupled with a pre-training paradigm that effectively mitigates codebook collapse and enhances bitrate efficiency.
@@
  • Extensive comparisons with Descript’s audio codec on a multilingual speech corpus demonstrate that transformers and CS-RVQ, the core components of ESC, are superior backbones for speech foundation models than the mainstream convolutions and RVQ.
@@ 2 Related Work
2.1 Neural Audio Codecs
Recently, most notable neural audio codecs have been based on the vector quantization (VQ) network, including SoundStream Zeghidour et al. (2021), EnCodec Défossez et al. (2023), and Descript’s audio codec (DAC) Kumar et al. (2023). SoundStream is distinguished as the first universal codec capable of handling diverse audio types. EnCodec improves compression rates by integrating a lightweight transformer language model within the discrete latent space and implements a streaming architecture. Building on similar backbones, Kumar et al. (2023) further explore the implications of quantization dropout, a technique for bitrate scalability, and demonstrate the superiority of periodic inductive bias functions over common activation functions for audio signal modeling gil Lee et al. (2023); Ziyin et al. (2020). These models directly process audio waveforms and are classified as time-domain codecs.
In contrast, frequency-domain codecs focus on processing more intuitive audio spectrogram features. Lyra Kleijn et al. (2021), for example, converts audio waveforms into log mel-spectrograms and directly quantizes them into tokens. Due to the non-invertibility of mel-spectrograms, it relies on a vocoder Kalchbrenner et al. (2018) for waveform synthesis. To circumvent the inefficiencies associated with heavy vocoders, some frequency-domain codecs, including TFNet Jiang et al. (2022b) and our ESC, employ the invertible Short-time Fourier Transform (STFT) to convert waveforms into complex spectra. This design enables the reconstructed STFT spectra to be seamlessly inverted back into waveforms without information loss using inverse-STFT. Among recent audio codecs, DAC achieves state-of-the-art compression ratios and reconstruction quality, though its computation bottlenecks are sometimes overlooked.
@@ 2.2 Swin Transformers
Vision Transformers (ViTs) Dosovitskiy et al. (2020) have outperformed convolutional neural networks (CNNs) in various image processing tasks, largely due to their superior ability to capture complex patterns. The Swin Transformer Liu et al. (2021), a notable variant, enhances this capability by employing a hierarchical approach with shifted window attention mechanisms, enabling it to scale efficiently to high-resolution signals while maintaining computational efficiency. In the context of image compression, Swin Transformers have demonstrated exceptional performance. Studies by Zhu et al. (2021) and Zou et al. (2022) show that Swin Transformers surpass CNNs in modeling spatial hierarchies and long-range dependencies. The attention mechanism facilitates the accurate preservation of essential details and textures, even at lower bitrates. These capabilities suggest that transformers could also be effective in applications beyond image compression, such as modeling audio spectrograms.
@@ 2.3 Vector Quantization
In the Vector Quantized Variational Autoencoder (VQ-VAE) (Van Den Oord et al., 2017), vector quantization (VQ) functions as a trainable layer that deterministically quantizes encoded latent variables by mapping them to their nearest neighbors in an embedding codebook. A VQ layer, denoted as $Q(\cdot\ ;\mathcal{C})$, is parameterized by a collection of continuous vectors $\mathcal{C}=\{\bm{c}_{1},...,\bm{c}_{K}\}$, each referred to as a codeword, with its associated index known as a code. The layer quantizes a vector $\bm{z}_{e}\in\mathbb{R}^{d}$ to $\bm{z}_{q}\in\mathbb{R}^{d}$ by selecting the Euclidean nearest codeword $\bm{c}_{k}$ from the codebook $\mathcal{C}$, i.e.,
@@
For convenience, we denote the output of the VQ function as $(\bm{z}_{q},\tilde{z}_{q})$, where $\tilde{z}_{q}:=k$ represents the discrete code corresponding to the nearest codeword. During compression, the encoding process outputs the discrete index $\tilde{z}_{q}$, which is stored with a $\log_{2}K$ bit budget. The decoding process starts by retrieving the continuous vector $\bm{z}_{q}$ from the codebook using the index $\tilde{z}_{q}$. The VQ function $Q(\cdot;\mathcal{C})$ is non-differentiable due to the $\operatorname*{arg\,min}$ operator. Common strategies use a straight-through estimator (STE) Bengio et al. (2013) to bypass this in back-propagation. In other words, the gradient component $\frac{\partial\bm{z}_{q}}{\partial\bm{z}e}$ is estimated by identity. Additionally, auxiliary losses including codebook loss and commitment loss are proposed to pull the codewords and latent features closer:
@@
Here $\text{sg}(\cdot)$ denotes the stop-gradient operator. The first term updates the codebook with an $l_{2}$ error, pushing the codewords towards the input vectors. The second term ensures that $\bm{z}_{e}$ commits to the embedding without growing arbitrarily. The scalar $\beta$ balances the importance of updating the codebook and the encoder.
@@ 2.4 Codebook Collapse
Straight-through estimators (STEs) can lead to significant issues, most notably codebook collapse, as detailed by Vuong et al. (2023). In a recent study, Huh et al. (2023) provide a plausible explanation, attributing the collapse to an internal codebook covariate shift during training. Frequent adjustments in encoder representations cause misalignment with the codebook, resulting in only a subset of codewords being updated. Consequently, VQ layers are prone to divergence, often ending up with a significant number of inactive vectors. Various strategies have been proposed in generative modeling context to address this issue, including stochastic quantization Takida et al. (2022); Zhang et al. (2023), self-annealed soft-to-hard quantization Agustsson et al. (2017), re-initializing codewords using K-means centroids every few epochs Łańcucki et al. (2020); Dhariwal et al. (2020), and reformulating with finite scalar quantization Mentzer et al. (2024). In audio compression, Kumar et al. (2023) address codebook collapse by down-projecting codewords Yu et al. (2022) and normalizing them within a Euclidean ball Łańcucki et al. (2020).
Figure 1: The framework of ESC: input speech is transformed to a complex STFT $\mathcal{X}$ and linearly embedded into patches. Encoder STBs iteratively halve the frequency resolution and produce hierarchical feature representations. Mirrored decoder STBs recover the frequency resolution by progressively leveraging coarse-to-fine quantized residual features between encoder and decoder hidden states. The entire network is solely composed of efficient transformer blocks and vector quantization layers. The figure displays a scenario when the deepest $3$ of $n+1$ total bitstreams (solid lines) are transmitted, with others left inactive.
@@ 3.1 Overall Architecture
As illustrated in Figure 1, ESC operates on the complex spectrum $\mathcal{X}\in\mathbb{R}^{2\times F\times T}$ derived from the Short-Time Fourier Transform (STFT) of a speech signal. Here, the real and imaginary components of $\mathcal{X}$ are treated as separate channels. Instead of using strided convolutions, ESC comprises a series of mirrored transformer encoder and decoder layers, each performing downsampling or upsampling to create coarse and fine representations, as described in Section 3.3. Starting from the quantized latents at the bottleneck VQ, the decoder progressively reconstructs the original spectrum by leveraging multi-level quantized residuals between the intermediate features of the encoder and decoder. This cross-scale decoding mechanism is further detailed in Section 3.4. Finally, the reconstructed spectrum $\hat{\mathcal{X}}$ is transformed back into a waveform through the inverse-STFT.
@@ 3.2 Notations
We first define some notations for clarity. The encoder and decoder are denoted by $F_{\phi}(\cdot)$ and $G_{\psi}(\cdot)$, respectively, each being a composition of individual layer functions $f_{\phi_{1}},\ldots,f_{\phi_{n}}$ and $g_{\psi_{1}},\ldots,g_{\psi_{n}}$. We use $\mathcal{Z}\in\mathbb{R}^{C\times F\times T}$ to denote a spectrum feature and $\bm{z}\in\mathbb{R}^{CF}$ to denote a flattened time frame vector in $\mathcal{Z}$. Specifically, $\mathcal{Z}_{e_{i}}$ refers to the feature after the $i$-th encoder layer, and $\mathcal{Z}_{q_{i}}$ denotes the $i$-th decoder feature.
@@
Here, $\mathcal{Z}_{e_{0}}$ is the original input feature and $\mathcal{Z}_{q_{0}}$ is the latent representation at the bottleneck.
@@ 3.3 Transformer Encoder and Decoder
Patchify. The encoder starts with a linear patchify module, where the complex spectrum $\mathcal{X}$ is divided into small patches and linearly up-projected:
@@
Here, the patch size across the frequency and temporal dimensions is $(\frac{F}{H_{0}},\frac{T}{W_{0}})$. This step reduces the input resolution to alleviate the computational burden on attention computation. At the end of the decoder, a symmetric de-patchify module reshapes the decoded patch feature $\mathcal{Z}_{q_{n}}$ and linearly down-projects it to produce a recovered spectrum $\hat{\mathcal{X}}$.
@@
Downsampling and upsampling. ESC maintains temporal resolution while scaling frequency resolution to equalize bitrates across different bitstreams. To achieve this, we modify the original patch merging/splitting modules with a single-dimensional pixel unshuffle/shuffle module Shi et al. (2016) along the frequency dimension. During encoder downsampling, an intermediate encoder spectrum feature $\mathcal{Z}_{e_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is first reshaped and then projected by $P_{e_{i}}\in\mathbb{R}^{vC_{i}\times C_{i+1}}$ as follows:
@@
where $v$ is the down-scaling factor. The upsampling process mirrors this operation in reverse. An intermediate decoder feature $\mathcal{Z}_{q_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is first projected by $P_{q_{i}}\in\mathbb{R}^{C_{i}\times vC_{i+1}}$ and then reshaped, resulting in an up-scaled frequency resolution:
@@
Overall, the transformer encoder and decoder layers are mirrored, creating symmetric and hierarchical representations of the input audio spectrum. With these backbones, ESC is a fully transformer-based codec without any convolutional modules.